OpenStudy (anonymous):

11. Add and simplify: (u+w/uw^2)+(2u+w/u^2w)

OpenStudy (anonymous):

\[\frac{u+w}{uw^2}+\frac{2u+w}{u^2w}=\frac{u^2+uw}{u^2w^2}+\frac{2uw+w^2}{u^2w^2}\] \[\frac{u^2+uw}{u^2w^2}+\frac{2uw+w^2}{u^2w^2}=\frac{u^2+3uw+w^2}{u^2w^2}\]

OpenStudy (anonymous):

this is simplified?